OpenStudy (anonymous):

(!!HELP, PLEASE!) How would I approach a related rates problem that asks me to find the rate of water leaking from an inverted conical tank? I was thinking dv/dt = dP/dt (the rate water is being added) - dL/dt (the unknown, or the rate of the leak). I know h - both of them. I'd have to find dv/dt, right? Please help!

OpenStudy (anonymous):

I was thinking of simply subtracting the derivatives once I find them.

OpenStudy (anonymous):

Chain rule:\[\frac{dy}{dx}=\frac{dy}{du}*\frac{du}{dx}\] Usually comes in handy when solving for related rates, but I actually don't know how to solve this question.

OpenStudy (precal):

http://www.mathscoop.com/calculus/derivatives/applications/related-rates-cylinder-problems.php try this website, it might help you

OpenStudy (anonymous):

@precal I'm still not understanding finding the rate of a leak....